# Quantum geometric formulation of Brans-Dicke theory for Bianchi I spacetime

### v2

@@ Acknowledgements.
This work was partly done when M.S. was visiting the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India. M.S. would like to express his sincere gratitude to Professor Dr. Naresh Dadhich for his kind invitation to visit IUCAA in the fall of 2022. The vibrant research environment and the scenic beauty of the campus was an impetus to carry this piece of work. L.L.G is supported by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307636/2023-2 and from the Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.297/2021. L.L.G. would like to thank IPMU - Kavli Institute for the Physics and Mathematics of the Universe for warm hospitality and also thank Prof. Elisa Ferreira for the constant scientific inspiring discussions. R.O.R. acknowledges financial support by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. A.W. is partially supported by the (US) NSF grant: PHY2308845.

### v3

@@ Acknowledgements.
This work was partly done when M.S. was visiting the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India. M.S. would like to express his sincere gratitude to Professor Dr. Naresh Dadhich for his kind invitation to visit IUCAA in the fall of 2022. The vibrant research environment and the scenic beauty of the campus was an impetus to carry this piece of work. L.L.G is supported by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307636/2023-2 and from the Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.297/2021. L.L.G. would like to thank IPMU–Kavli Institute for the Physics and Mathematics of the Universe for warm hospitality and also thank Professor Elisa Ferreira for the constant scientific inspiring discussions. R.O.R. acknowledges financial support by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. A.W. is partially supported by the US NSF Grant No. PHY2308845.